Compared to HCO,, HCO, has a A) Higher ka B) Lower Ka C) Lower pka D) Higher pka E) Both (A) and (C)
Answer #1

H2CO3 is stronger acid than HCO3- .Ka values are given below.

For H2CO3  , Ka = 4.5×10-7 , pKa = 6.35

Ka of HCO3- = 4.7×10-11 and pKa = 10.33.

[ pKa = - logKa]

So, compare to HCO3- , H2CO3 has higher Ka and lower pKa .

Option A and C are correct.